OpenStudy (anonymous):

Algebra 2 Help please!!! Ill give a medal and become a fan :D

OpenStudy (anonymous):

ill help

OpenStudy (anonymous):

The half life of a certain radioactive material is 36 days. An initial amount of the material has a mass of 487 kg. Write an exponential function that models the decay of this material. Find how much radioactive material remains after 5 days. Round your answer to the nearest thousandth. a)y=487(1/2)^36x ; 0.318kg b)y=2(1/487)^1/36x ; 0.847kg c)y=487(1/2)^1/36x ; 442.302kg d)y=2(1/487)^36x ; 0kg

OpenStudy (anonymous):

@joseph211877

OpenStudy (anonymous):

Thats the problem^

OpenStudy (anonymous):

A

OpenStudy (anonymous):

@joseph211877 Thanks :)

OpenStudy (anonymous):

your very well come :)

OpenStudy (anonymous):

I think its C but that's just me.

OpenStudy (anonymous):

A = A(o)[0.5^(t/t(1/2)] A = 487[0.5^5/36] = 487(0.9082) = 442.302kg

OpenStudy (anonymous):

Looking at the initial and final amount, initial amount of 487, and final amount of 442.3. The half life is 36 days and only 5 days have passed, meaning we need to have more than 243.5 which only C has.

OpenStudy (anonymous):

yah hes right its C sorry

OpenStudy (anonymous):

After the semicolon, is the amount left. 36x, means we multiply each day by 36, instead of (1/36)x, which means we multiple each day by 1/36, ty ty.

OpenStudy (anonymous):

Thanks guys! I changed it:) @christopher_miller420 @joseph211877

OpenStudy (anonymous):

yup just doing my job :P